An elevator accelerates upward at 1.2 m/s2.
The acceleration of gravity is 9.8 m/s2. What is the upward force exerted by the
floor of the elevator on a(n) 67 kg passenger?
Answer in units of N.
Don't round answer.
1
Expert's answer
2019-10-16T09:27:32-0400

Using the second Newton's law:


"ma=F-mg"

The upward force exerted by the floor of the elevator on a(n) 67 kg passenger

"F=m(a+g)=67(1.2+9.8)=737\ N"